# Hydrodynamic noise in one dimension: projected Kubo formula and its vanishing in integrable models

### v1

@@
and its vanishing in integrable models
@@
Hydrodynamic noise is the Gaussian process that emerges at larges scales of space and time in many-body systems. It arises by the central limit theorem applied to local microcanonical averages, representing the degrees of freedom that have been forgotten when projecting coarse-grained observables onto conserved quantities. It comes with “bare” diffusion terms. In one dimension of space, nonlinearities of the hydrodynamic equation are relevant (from a renormalisation perspective), usually giving rise to hydrodynamic superdiffusion. But in linearly degenerate systems, where the relevant nonlinearity vanishes, the diffusive scaling stays intact. Nevertheless, anomalies remain. We show that in such systems, the noise covariance is determined in terms of a modification of the Kubo formula, where effects of ballistic long-range correlations have been subtracted. This is the projected Onsager matrix, in which so-called quadratic charges are projected out. We show that the Einstein relation holds, giving a projected bare diffusion, and that the remaining nonlinearities are tamed by a point-splitting regularisation. Putting these ingredients together, we obtain an exact and well-defined hydrodynamic fluctuation theory in the ballistic scaling of space-time, for the asymptotic expansion in the inverse variation scale, including the first subleading (diffusive-scale) corrections beyond large deviations. This is expressed as a stochastic PDE. We then obtain the anomalous hydrodynamic equation, which takes into account separately long-range correlations and bare diffusion. Using these result, in integrable systems, we show that hydrodynamic noise must be absent, as was conjectured recently.
@@ 1 Introduction
At large scales of space and time, degrees of freedom in many-body systems separate into “fast” and “slow” ones. Fast ones can be assumed to quickly relax throughout space-time, at each time taking values that are determined by the spatial profile of slow ones. This is the projection principle of Mori and Zwanzig [1, 2], and slow degrees of freedom are associated with conserved quantities [3, 4, 5, 6]. At the linearised level, it is called the Boltzmann-Gibbs principle, and says that fluctuation fields for arbitrary observables linearly project onto fluctuation fields for conserved densities [6].
@@
More precisely, on short times we may assume coarse-grained conserved densities to be fixed, while generic coarse-grained observables fluctuate a lot, thus converging to their microcanonical values, functions of the conserved densities, in each “fluid cell”. This is exact in the strict limit where spatial variations (variations of particle or energy densities, etc.) are on large wavelengths $\ell$, as fluid cells can be taken as large as we want. But looking at correction terms in the hydrodynamic expansion in $1/\ell$, this projection is not perfect. The microcanonical averages on fluid cells receive many corrections. In particular, by the central limit theorem, there must be an emergent Gaussian white noise associated to each coarse-grained observable. This noise is delta-correlated in space-time, but correlated amongst different observables. Its covariance is usually assumed to be given by the Kubo formula for the Onsager matrix, and by the fluctuation-dissipation theorem, this noise must be associated with dissipation, giving viscous and diffusive terms by the Einstein relation [4, 5, 6, 7]. This is the basis for the theory of fluctuating hydrodynamics. It is valid for both deterministic and stochastic systems. In particular, in the latter, the emergent noise is in general not simply reIated to the microscopic noise. In purely diffusive systems, where fluxes vanish, and under the diffusive scaling of space and time $\vec{{\rm x}}\propto\ell,\,\mathrm{t}\propto\ell^{2}$, the hydrodynamic expansion in fact starts at the diffusive level, and this is the basis for the Macroscopic Fluctuation Theory (MFT) [8], a large-deviation theory.
In one dimension of space, if fluxes are non-zero, such as in Galilean or relativistic gas or in integrable models, their nonlinearity as functions of conserved densities are generically relevant in the sense of the renormalisation group, and leads to superdiffusion. For instance, take a system with a single conservation law $\partial_{\mathrm{t}}q+\partial_{\mathrm{x}}j=0$. One writes the coarse-grained current $\overline{j}$ as a function of the coarse-grained conserved density $\overline{q}$
$\overline{j}=\mathsf{j}(\overline{q})-\frac{1}{2}\mathfrak{D}\partial_{\mathrm{x}}\overline{q}+\xi$ (1)
where $\mathsf{j}(\overline{q})$ is the microcanonical average – the flux –, and $\xi$ is the noise, $\langle\!\langle\xi(\mathrm{x},\mathrm{t})\xi(\mathrm{x}^{\prime},\mathrm{t}^{\prime})\rangle\!\rangle=\mathfrak{L}\delta(\mathrm{x}-\mathrm{x}^{\prime})\delta(\mathrm{t}-\mathrm{t}^{\prime})$, with the associated diffusion function $\mathfrak{D}=\mathfrak{L}/\chi$ where $\chi$ is the susceptibility. The fluctuating hydrodynamics is the stochastic PDE
$\partial_{\mathrm{t}}\overline{q}+\partial_{\mathrm{x}}\Big{(}\mathsf{j}(\overline{q})-\frac{1}{2}\mathfrak{D}\partial_{\mathrm{x}}\overline{q}+\xi\Big{)}=0.$ (2)
This equation is generically singular and not well-defined. However, it is known that if the second derivative $\mathsf{j}^{\prime\prime}$ is non-vanishing, then the result is in the KPZ universality class under the appropriate scaling of space and time, giving superdiffusion. With more conservation laws, $\partial_{\mathrm{t}}q_{i}+\partial_{\mathrm{x}}j_{i}=0$, $i=1,2,\ldots$, similar effects arise if certain coefficients, combinations of flux derivatives, do not vanish, see [9]. These coefficients can be written as [10]
$\frac{\partial^{2}\mathsf{j}_{i}}{\partial\overline{q_{I}}^{2}}$ (3)
where densities $q_{I}$ are in the “normal mode” basis (see Appendix A).
@@
But there is a class of models where superdiffusion does not arise: linearly degenerate systems. Linear degeneracy means that hydrodynamic modes do not self-interact: the hydrodynamic velocity of any normal mode, does not depend on this normal mode [11, 12, 13] (see also [14]). These are still interacting, yet close to free systems. They do not develop shocks [15, 16, 17, 18], and satisfy (3) [19], amongst other properties. Integrable many-body systems have hydrodynamic equations that are linearly degenerate, and there are many other examples [20, 21, 22, 23]. Because there is no shock, the fluctuating theory at the leading order in the ballistic scaling
${\rm x}=\ell x,\quad\mathrm{t}=\ell t,\quad\ell\to\infty$ (4)
is simply the deterministic transport of initial fluctuations via the Euler equation. This is the Ballistic Macroscopic Fluctuation Theory (BMFT) [24, 25], a large-deviation theory paralleling the MFT. But beyond this, for the $1/\ell$ corrections to the large-deviation behaviour, the many-component equivalent of the fluctuating hydrodynamics (2) still generically is nonlinear. Further, it was recently conjectured that in integrable systems, there is no noise, $\mathfrak{D}=\xi=0$ [20, 21, 19], and this was proposed to hold in linearly degenerate systems more generally [19].
@@
In this paper, we propose the exact fluctuating hydrodynamic theory for linearly degenerate systems, in the ballistic scaling (4), up to, including, the first subleading $1/\ell$ order of the asymptotic expansion. The resulting stochastic PDE, Eq. (25) below, is well defined for the asymptotic expansion in $1/\ell$. Along with the nonlinear, fluctuating Boltzmann-Gibbs principle (3.1) for generic observables, it allows one to evaluate all connected correlation functions as $\ell\to\infty$, both their leading BMFT large-deviation form, worked out in [25, 26], and the first subleading order in $1/\ell$. Likewise, a field-theoretic reformulation of the stochastic PDE, such as that proposed in [25, Sec 3.3], should provide large deviations and their first subleading corrections. These will be studied in a future work.
The crucial points of this paper are as follows. (1) We show that the fluxes are subject to a “regularisation”, coming from the non-fluctuating, non-diffusive correction to the fluid-cell microcanonical average, in inverse power of the cell’s size. This regularisation, in turn, guarantees that the stochastic PDE is well defined in the $1/\ell$ expansion, despite delta-correlations both in noise and initial conditions. It is the point-splitting regularisation first conjectured in [19]. (2) We show that the noise covariance is not given by the Kubo formula for the Onsager matrix, but by a “projected” version of it. In this version, quadratic charges [10] are projected out. These represent the effects of long-range correlations [24], coming from the nonlinearity of the flux, on the diffusive scale, as worked out in [19] in the context of integrability. The fact that flux nonlinearities give rise to diffusive contributions was first proposed as “diffusion from convection” in [27] in the context of integrability, and our proof uses a calculation that is somewhat similar. (3) We show that the diffusion matrix is connected to the projected Onsager matrix by the Einstein relation. (4) We show that in integrable models, the noise for currents vanishes. This follows from our result, along with earlier calculations of the full Onsager matrix in integrable models [28, 4, 29, 30, 31]. We also derive it from first principles using the infinite family of commuting flows that many-body integrable models admit. This shows the conjecture of [20, 21, 19]. We emphasise, however, that in generic linearly degenerate systems, the noise does not vanish – this is a property of integrability.
The projected Kubo formula for the noise is also worked out, albeit in the diffusive instead of ballistic scaling, in the simultaneous paper [23] of which we have been made aware. In particular, this is evaluated in a specific model, where it is shown to agree with microscopic calculations and to be non-zero.
We emphasise that bare diffusion is different from hydrodynamic diffusion. Linearised hydrodynamics for two-point correlatoin functions is diffusive. There, however, the hydrodynamic diffusion constant is connected by the Einstein relation to the full Onsager matrix, not the projected one (see e.g. [7]). Thus, even integrable models have diffusive two-point functions. Beyond the linearised level, however, these diffusive effects desintegrate into long-range correlations [19].
@@
From our fluctuating hydrodynamic theory, write down the hydrodynamic equation up to, including, the diffusive order $1/\ell$, see Eqs. (48), (49) below. It is a simple modification of that proposed in [19] for integrable models. Long-range corelations and bare diffusion are explicitly taken into account.
@@
We believe our derivation gives a better understanding of the microscopic origin of hydrodynamic noise and bare diffusion, and of the subtle principles behind fluctuating hydrodynamics more generally.
@@
The paper is organised as follows. In Section (2) we specify our general setup, describing the conserved quantities and states. In Section 3, we express our main result for the fluctuating hydrodynamic theory, and provide arguments as to how coarse-graining give rise to the special effects seen beyond the Euler scale, including the noise. In Section 4, we prove points (1)-(3) described above and write down the hydrodynamic equation. In Section 5 we show the absence of noise at the diffusive level in integrable systems. Finally, in Section 6 we conclude.
@@ 2 Conservation laws and states
Consider a dynamical system in one spatial dimension, which may be deterministic or stochastic, classical or quantum, systems with short-range, translation invariant interactions. Local observables $o({\rm x},\mathrm{t})$ are functions on phase space, random variables of stochastic processes, or operators on a Hilbert space, etc, which are supported on finite regions in microscopic units, around position ${\rm x}\in{\mathbb{R}}$ at time $\mathrm{t}\in{\mathbb{R}}$. We will write ${\rm x}=\ell x,\,\mathrm{t}=\ell t$ for the space-time point coordinates $(x,t)$ in macroscopic unit, where $\ell$ is the large macroscopic scale.
@@
where $q_{i}({\rm x},\mathrm{t})$ is the associated local density. The index $i$ lies in some index set which may be finite or not. The index may even be continuous like in integrable models where it represents the spectral space (such as velocities of quasiparticles); we still use the notation $\sum_{i}$. We assume that there are associated local currents $j_{i}({\rm x},\mathrm{t})$ giving rise to local conservation laws:
$\partial_{t}q_{i}({\rm x},\mathrm{t})+\partial_{\rm x}j_{i}({\rm x},\mathrm{t})=0.$ (6)
When discussing hydrodynamics at the diffusive order, the concept of PT symmetry plays an important role, as first discussed in [30]. We assume that there is an involution ${\mathcal{P}T}$ of the algebra of observables that reverse space and time, and that preserves all conserved densities and their currents. Further, we will only consider local observables $o$ that are real or hemitian, and that are preserved by the PT transformation. That is,
@@
This simplifies the discussion in subtle ways (we will use such PT invariance at one crucial point). Note that in general, there may be observables $o$ with negative eigenvalue under the ${\mathcal{P}T}$ involution; however, our results do not apply to such observables.
The macrocanonical ensemble is that which maximises entropy states under the constraints of the extensive conserved quantities. Formally, it has the Gibbs form
@@
where $\underline{\beta}=(\beta^{1},\beta^{2},\ldots)$ are the “Lagrange parameters”, the charges $Q_{i}=\int_{\mathcal{V}}{\rm d}^{d}{\rm x}\,q_{i}(\bm{{\rm x}})$ lie on the volume $\mathcal{V}$, and the large-volume limit $\mathcal{V}\to{\mathbb{R}}$ is assumed to exist for averages of local observables. $\mu$ is a time-invariant, homogeneous prior measure, for instance in classical Hamiltonian systems with canonical coordinates $({\rm x}_{a},{\rm p}_{a})$ one chooses ${\rm d}\mu=\sum_{N=0}^{\infty}(N!)^{-1}\prod_{a=1}^{N}{\rm d}{\rm x}_{a}{\rm d}{\rm p}_{a}$ (direct sum); while in quantum systems, $\langle\cdots\rangle_{\underline{\beta}}=\frac{{\rm Tr}\,{\rm e}^{-\sum_{i}\beta^{i}Q_{i}}\cdots}{{\rm Tr}\,{\rm e}^{-\sum_{i}\beta^{i}Q_{i}}}$. We assume that $\beta^{i}$’s are away from phase transitions, so that correlations are short-range in these states, and averages of local observables are smooth in $\underline{\beta}$. Likewise, the microcanonical ensemble is denoted (by a slight abuse of notation)
$\langle\cdots\rangle_{\underline{\mathsf{q}}}=\lim_{\mathcal{V}\to{\mathbb{R}}^{d}}\frac{\int_{Q_{i}/|\mathcal{V}|\,\in\,[\mathsf{q}_{i}-\epsilon,\mathsf{q}_{i}+\epsilon]\,\forall i}{\rm d}\mu\,\cdots}{\int_{Q_{i}/|\mathcal{V}|\,\in\,[\mathsf{q}_{i}-\epsilon,\mathsf{q}_{i}+\epsilon]\,\forall i}{\rm d}\mu}$ (9)
@@
is assumed to be positive,
@@
so that the map $\underline{\beta}\mapsto\langle\underline{q}\rangle_{\underline{\beta}}$ is invertible. This defines functions $\beta^{i}(\underline{\mathsf{q}})=\beta^{i}(\mathsf{q}_{1},\mathsf{q}_{2},\ldots)$ for all $i$’s. There is equivalence between microcanonical and macrocanonical ensembles [32, 33, 34]: for every local obsevable $o$, we have $\langle o\rangle_{\underline{\mathsf{q}}}=\langle o\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$. Throughout, we denote the resulting average of $o$, as a function of conserved densities $\underline{\mathsf{q}}$, by using the “sans-serif” font for the observable,
@@
for a large, macroscopic scale $\ell$. This state has short range correlations. More general states, with long-range correlations or with long-wavelength source insertions at various macroscopic times, can also be considered.

### v2

# Hydrodynamic noise in one dimension: projected Kubo formula and how it vanishes in integrable models

@@
and how it vanishes in integrable models
@@
Hydrodynamic noise is the Gaussian process that emerges at larges scales of space and time in many-body systems. It is justified by the central limit theorem, and represents degrees of freedom forgotten when projecting coarse-grained observables onto conserved quantities. It is the basis for fluctuating hydrodynamics, where it appears along with “bare” diffusion terms related to the noise covariance by the Einstein relation. In one dimension of space, nonlinearities are relevant and may modify the corrections to ballistic behaviours by superdiffusive effects. But in systems where no shocks appear, such as linearly degenerate systems and integrable systems, the diffusive scaling of these corrections stays intact. Nevertheless, anomalies remain. We show that in such systems, the noise covariance is given by a modification of the Kubo formula, where effects of ballistic long-range correlations have been projected out. We further show that nonlinearities are tamed by a point-splitting regularisation. With these ingredients, we obtain a well-defined hydrodynamic fluctuation theory in the ballistic scaling of space-time, as a stochastic PDE. It describes the asymptotic expansion in the inverse variation scale of connected correlation functions, self-consistently organised via a cumulant expansion. The resulting anomalous hydrodynamic equation, for average conserved densities, takes into account both long-range correlations and bare diffusion, generalising a recent work by Hübner, Biagetti, De Nardis, and Doyon. Despite these anomalies, two-point functions satisfy an ordinary diffusion equation, with diffusion matrix determined by the Kubo formula. In integrable systems, we show that hydrodynamic noise, hence bare diffusion, must vanish, as was conjectured recently, and argue that under an appropriate gauge of the currents, this is true at all orders. Thus the Ballistic Macroscopic Fluctuation Theory give the all-order hydrodynamic theory for integrable models.
@@ 1 Introduction
At large scales of space and time, degrees of freedom in many-body systems separate into “fast” and “slow” ones. Fast ones can be assumed to quickly relax throughout space-time, at each time taking values that are determined by the spatial profile of slow ones. This is the projection principle of Mori and Zwanzig [1, 2]. Slow degrees of freedom are associated with conserved quantities [3, 4, 5, 6]: coarse-grained conserved densities vary slowly due to interchanges at the surface of “fluid cells”, while generic coarse-grained observables fluctuate a lot, thus converging to their microcanonical values in each fluid cell. At the linearised level, this is called the Boltzmann-Gibbs principle [6], and keeping microcanonical values as nonlinear functions of densities, it is the basis for the Ballistic Macrocopic Fluctution Theory (BMFT) [7, 8], a large-deviation theory.
@@
This is exact in the strict limit where spatial variations are on infinite length scales $\ell$, as fluid cells can be taken as large as we want. But looking at corrections in the hydrodynamic expansion in $1/\ell$, this projection is not perfect. The microcanonical averages on fluid cells receive many corrections. In particular, by the central limit theorem, there must be an emergent Gaussian white noise for each coarse-grained observable. This noise is usually assumed to be delta-correlated in space-time, but correlated amongst different observables, and by the fluctuation-dissipation theorem and Einstein relation, is argued to be associated with diffusion terms [4, 5, 6, 9]. This is the basis for the theory of fluctuating hydrodynamics, a stochastic PDE. It applies both to deterministic systems, where the noise is an emergent phenomenon carrying initial fluctuations, and stochastic systems, where the emergent hydrodynamic noise is in general not simply related to the microscopic noise. Its applicability to many-body quantum systems has also been investigated [10].
In driven-diffusive systems, and under the diffusive scaling of space and time $\vec{{\rm x}}-\vec{v}\mathrm{t}\propto\ell,\,\mathrm{t}\propto\ell^{2}$, the combination of fluctuating hydrodynamics and fluctuations of initial conditions gives rise to the Macroscopic Fluctuation Theory (MFT) [11], a large-deviation theory.
But fluctuating hydrodynamics goes beyond, and holds also when ballistic fluxes – microcanonical averages of currents – are nonlinear functions. For instance, take a system with a single conservation law $\partial_{\mathrm{t}}q+\partial_{\mathrm{x}}j=0$, where $\mathrm{x},\mathrm{t}$ are microscopic coordinates. One writes the coarse-grained current $\overline{j}$ as a function of the coarse-grained conserved density $\overline{q}$
$\overline{j}=\mathsf{j}(\overline{q})-\frac{1}{2}\tilde{\mathfrak{D}}(\overline{q})\partial_{\mathrm{x}}\overline{q}+\eta$ (1)
where $\mathsf{j}(\overline{q})$ is the microcanonical average of the current in a state characterised by $\overline{q}$, and $\eta({\rm x},\mathrm{t})$ is an emergent noise, correlated on microscopic space-time scales only, with strength $\tilde{\mathfrak{L}}(\overline{q})=\tilde{\mathfrak{D}}(\overline{q})\chi(\overline{q})$ where $\chi(\overline{q})$ is the susceptibility. The fluctuating hydrodynamics, valid on large scales of $\mathrm{x},\mathrm{t}$, is the stochastic PDE
$\partial_{\mathrm{t}}\overline{q}+\partial_{\mathrm{x}}\Big{(}\mathsf{j}(\overline{q})-\frac{1}{2}\tilde{\mathfrak{D}}(\overline{q})\partial_{\mathrm{x}}\overline{q}+\eta\Big{)}=0.$ (2)
Crucially, the noise strength $\tilde{\mathfrak{L}}$ and bare diffusion $\tilde{\mathfrak{D}}$ are generically different from their macroscopic counterparts [12, 13, 14, 15, 16]. The evaluation of bare transport coefficients is delicate [13, 17, 18, 19, 20], and the understanding of the effect of hydrodynamic noise on large-scale physics is an important question, potentially related to spontaneous stochasticity [21] and the dissipation anomaly in turbulence [22]. For instance, in low-dimensional systems, $\tilde{\mathfrak{L}}$ is not given by the Onsager coefficient – the Green-Kubo formula from the current-current correlator. In fact, the Onsager coefficient often diverges on infinite volumes, indicating the presence of superdiffusion. Notably, if the second derivative $\mathsf{j}^{\prime\prime}$ is non-vanishing, Eq. (2), under the KPZ scaling of space and time, is known to lie in the KPZ universality class. With more conservation laws, $\partial_{\mathrm{t}}q_{i}+\partial_{\mathrm{x}}j_{i}=0$, $i=1,2,\ldots$, the KPZ universality class also emerges [23], as controlled by the “diagonal 3-point couplings”, equilibrium 3-point correlation functions $\mathsf{A}_{k}^{~II}=\langle Q_{I},Q_{I},j_{k}^{-}\rangle^{\rm c}$ simply related to the ballistic flux derivatives [24] (see Appendix A).
@@
Interestingly, one can show that if the Euler-scale equation is such that shocks are not produced, then the diagonal 3-point couplings vanish, $\mathsf{A}_{k}^{~II}=0$, thus there is no superdiffusion, see [25] and Appendix C.1. A class of such no-shock systems is that of linearly degenerate systems. Linear degeneracy was first introduce by Lax [26], see also [27, 28, 29] and [30, Sec 3.2]. It means that hydrodynamic modes do not self-interact: the hydrodynamic velocity of any normal mode, does not depend on this normal mode, which precludes the appearance of shocks [31, 32, 30]. Integrable many-body systems have linearly degenerate hydrodynamic equations [28, 29, 33] and are proven not to develop shocks [34, 35], and there are many other examples [36, 37, 38, 39].
But despite the absence of superdiffusion, there are still large-scale anomalies in linearly degenerate systems.
Consider the the ballistic scaling
${\rm x}=\ell x,\quad\mathrm{t}=\ell t,\quad\ell\to\infty.$ (3)
In no-shock systems, the fluctuating theory at the leading order is simply the deterministic transport of initial fluctuations via the Euler equation: this makes sense as the absence of shocks makes solutions unique, and entropic effects do not occur. This is the BMFT [7, 8]. But beyond the leading order, noise and diffusion appears: under the scaling (3) we may expect to have non-zero bare contributions $\frac{1}{2\ell}\tilde{\mathfrak{D}}(\overline{q})\partial_{x}\overline{q}+\frac{1}{\ell}\xi$ in (2), where $\xi$ is delta-correlated in space-time. The question of the exact diffusive scale fluctuating hydrodynamics, bare transport coefficients, and their effects on large-scale correlation functions and the hydrodynamic equation for average densities, has not been addressed. In fact, it was recently conjectured that in integrable systems, there is no bare diffusion neither noise, $\xi=0$ [36, 37, 25], even though there are diffusive-scale effects on the hydrodynamic equation [25], and two-point correlation functions satisfy a diffusive equation [40, 41, 42].
@@
In this paper, we propose the exact diffusive scale fluctuating hydrodynamic theory for no-shock systems in the ballistic scaling (3). Along with a nonlinear fluctuating Boltzmann-Gibbs principle for generic observables, Eq. (21) below, this gives a framework for evaluating the leading and first subleading order in the asymptotic $\ell\to\infty$ expansion of all $n$-point connected correlation functions of local observables, $n=1,2,3,\ldots$. The leading order is given by the BMFT and worked out in [8, 43]. The first subleading order is a $1/\ell$ relative correction. The resulting stochastic PDE, Eq. (26) below, is well defined for this asymptotic expansion. In particular, we evaluate exactly the noise strength and bare diffusion in terms of a modification of the Green-Kubo formula. Our findings are as follows:
(1) We show that there is a correction to the fluid-cell microcanonical average, in inverse power of the cell’s size. This correction is usually not discussed in works on fluctuating hydrodynamics, but is crucial. It means that fluxes are subject to a “regularisation”, $\mathsf{j}(\overline{q})\to\mathsf{j}^{\rm reg}(\overline{q})$, which we show is the point-splitting regularisation first conjectured in [25]. This in turn guarantees that the stochastic PDE is well defined in the $1/\ell$ expansion, despite delta-correlations (in macroscopic coordinates) in initial conditions.
(2) We show that the noise covariance is given by a “projected” version of the Green-Kubo formula. In this version, quadratic charges [24] are projected out. These represent the effects of wave interaction which are at the source of long-range correlations [7], and come from the nonlinearity of the flux as worked out in [25] in the context of integrability. The fact that flux nonlinearities give rise to diffusive contributions was first proposed as “diffusion from convection” in [44] in the context of integrability, and our proof uses a calculation that is somewhat similar.
(3) We show that the bare diffusion matrix is connected to the projected Onsager matrix by the Einstein relation. This relation is natural on the grounds of the fluctuation-dissipation theorem, but it remains an assumption in the context of fluctuating hydrodynamics. We show that it is a consequence of the self-consistency of the PDE as a principle for the $1/\ell$ expansion, along with the conservation laws.
(4) We show that in many-body integrable models, the noise for currents vanishes. This follows from our result, along with earlier calculations of the full Onsager matrix in integrable models [45, 5, 40, 41, 42] (that used different methods). This establishes the conjecture of [36, 37, 25]. This implies that in integrable systems, initial fluctuations do not influence the diffusive corrections to coarse-grained nonlinear currents as functions of conserved densities. We also propose a more heuristic derivation of the vanishing of noise from first principles, using the infinite family of commuting flows that many-body integrable models admit. The argument shows that under an appropriate choice of “gauge” for the currents, their noise vanish at all orders in $1/\ell$. It also effectively gives a new, first-principle derivation of the exact formula for the full Onsager matrix in integrable models. We emphasise, however, that in generic linearly degenerate systems, the noise does not vanish (contrary to the conjecture made in [25]) – this is a property of integrability.
@@
From our fluctuating hydrodynamic theory, we write down the hydrodynamic equation up to, including, the diffusive order $1/\ell$, see Eqs. (LABEL:hydro1), (LABEL:hydro2) below. It is a simple modification of that proposed in [25] for integrable models. Long-range corelations and bare diffusion are explicitly taken into account separately.
Interestingly, the projected Kubo formula for the noise is also worked out, but in the diffusive instead of ballistic scaling, in the simultaneous paper [39] of which we have been made aware at the time of writing. Thus this formula holds no matter the choice of scaling. In particular, in [39] it is evaluated in a specific model, where it is shown to agree with microscopic calculations and to be non-zero.
We emphasise that bare diffusion is different from hydrodynamic diffusion. Linearised hydrodynamics for two-point correlation functions in linearly degenerate systems is not anomalous, and simply diffusive. There, the hydrodynamic diffusion constant is connected by the Einstein relation to the full Onsager matrix, not the projected one (see e.g. [9]). We show that this follows from fluctuating hydrodynamics theory; it was also argued for in integrable models in [25] using different ideas, and assumed to be true for some time [40, 41, 42]. Thus, integrable models have normal diffusive corrections to ballistic two-point functions. Beyond the linearised level, however, these diffusive effects desintegrate into long-range correlations [25]; higher-point functions have anomalous diffusive-scale corrections.
@@
We believe our general explanations and precise derivation give a better understanding of the microscopic origin of hydrodynamic noise and bare diffusion, and of the subtle principles behind fluctuating hydrodynamics more generally.
@@
The paper is organised as follows. In Section 2 we specify our general setup, describing the conserved quantities and states. In Section 3, we express our main result for the fluctuating hydrodynamic theory, and provide arguments as to how coarse-graining gives rise to the special effects seen beyond the Euler scale, including the noise. In Section 4, we prove points (1)-(3) described above. In Section 5 we show the absence of noise, point (4) above, in integrable systems. In Section 6, we obtain, from the fluctuating theory, the hydrodynamic equation for one-point averages out of equilibrium, and for two-point correlations in stationary states. Finally, in Section 7 we conclude.
Note added: Shortly after the first version of this paper was posted on arXiv, a paper by F. Hübner was also posted [46] which presents an analytical and numerical study of the hard rods model confirming our finding that hydrodymamic noise in integrable systems vanish at the diffusive level.
@@ 2 Conservation laws and states
Consider a dynamical system in one spatial dimension, which may be deterministic or stochastic, with short-range, translation invariant interactions. Fluctuating hydrodynamics, where a classical noise is assumed to emerge, is naturally a classical theory, and quantum effects may affect this in quantum systems. Many of the results below may still be applicable to quantum models [10], however we will not discuss this.
In order to have a universal description, one powerful method is to concentrate on local observables – such as in the algebraic formulation of statistical mechanics [47, 48]. Roughly, for our purposes, a local observable, denoted $o({\rm x},\mathrm{t})$, is a function on phase space, or a random variable of the stochastic process, etc, which is supported (in a natural sense) on some finite region of space in microscopic units, around position $\rm x$$\in{\mathbb{R}}$ at time $\mathrm{t}\in{\mathbb{R}}$. Observables can be multiplied with each other (so they form an algebra), and states (see below) are positive linear maps on this algebra.
We will write ${\rm x}=\ell x,\,\mathrm{t}=\ell t$, defining the space-time point coordinates $(x,t)$ in macroscopic unit, where $\ell$ is the large macroscopic scale [^1] We consistently use the different fonts ${\rm x},\mathrm{t}$ and $x,t$ for microscopic and macroscopic coordinates, respectively..
@@
where $q_{i}({\rm x},\mathrm{t})$ is the associated local density. The index $i$ lies in some index set $\mathcal{I}$ which may be finite or not. The index may even be continuous like in integrable models where it represents the spectral space (such as velocities of quasiparticles); we still use the notation $\sum_{i}$. We assume that there are associated local currents $j_{i}({\rm x},\mathrm{t})$ giving rise to local conservation laws:
$\partial_{\mathrm{t}}q_{i}({\rm x},\mathrm{t})+\partial_{\rm x}j_{i}({\rm x},\mathrm{t})=0.$ (5)
Total charges $Q_{i}$, on infinite volume, do not make sense as observables because they are infinite, but make sense within connected correlation functions, which is how we will use them.
When discussing hydrodynamics at the diffusive order, the concept of PT symmetry plays an important role, as first discussed in [41]. We assume that there is an involution $\mathcal{P}T$} of the algebra of observables that reverses space and time, and that preserves all conserved densities and their currents. Further, we will only consider local observables $o$ that are real or hemitian, and that are preserved by the PT transformation. That is,
@@
This simplifies the discussion in subtle ways. Note that in general, there may be observables $o$ with negative eigenvalue under the $\mathcal{P}T$} involution; however, our results do not apply to such observables.
The macrocanonical ensemble is that which maximises entropy under the constraints of the extensive conserved quantities. Formally, it has the Gibbs form
@@
where $\underline{\beta}=(\beta^{1},\beta^{2},\ldots)$ are the “Lagrange parameters”, the charges $Q_{i}=\int_{\mathcal{V}}{\rm d}^{d}{\rm x}\,q_{i}(\bm{{\rm x}})$ lie on the volume $\mathcal{V}$, and the large-volume limit $\mathcal{V}\to{\mathbb{R}}$ is assumed to exist for averages of local observables. $\mu$ is a time-invariant, homogeneous prior measure, for instance in classical Hamiltonian systems with canonical coordinates $({\rm x}_{a},{\rm p}_{a})$ one chooses ${\rm d}\mu=\sum_{N=0}^{\infty}(N!)^{-1}\prod_{a=1}^{N}{\rm d}{\rm x}_{a}{\rm d}{\rm p}_{a}$ (direct sum). We assume that $\beta^{i}$’s are away from phase transitions, so that correlations are short-range in these states, and averages of local observables are smooth in $\underline{\beta}$. Likewise, the microcanonical ensemble is denoted (by a slight abuse of notation)
$\langle\cdots\rangle_{\underline{\mathsf{q}}}=\lim_{\mathcal{V}\to{\mathbb{R}}}\frac{\int_{Q_{i}/|\mathcal{V}|\,\in\,[\mathsf{q}_{i}-\epsilon,\mathsf{q}_{i}+\epsilon]\,\forall i}{\rm d}\mu\,\cdots}{\int_{Q_{i}/|\mathcal{V}|\,\in\,[\mathsf{q}_{i}-\epsilon,\mathsf{q}_{i}+\epsilon]\,\forall i}{\rm d}\mu}$ (8)
@@
is always non-negative, and must be assumed to be positive for conserved quantities to have a hydrodynamic meaning [^3] This guarantees convexity of the free energy, a fundamental concept of statistical mechanics. See e.g. [51] for a general theorem concerning this in the context of the linearised Euler equation.. That is,
@@
and the map $\underline{\beta}\mapsto\langle\underline{q}\rangle_{\underline{\beta}}$ is invertible. This defines functions $\beta^{i}(\underline{\mathsf{q}})=\beta^{i}(\mathsf{q}_{1},\mathsf{q}_{2},\ldots)$ for all $i$’s. There is equivalence between microcanonical and macrocanonical ensembles [49, 50, 52]: for every local obsevable $o$, we have $\langle o\rangle_{\underline{\mathsf{q}}}=\langle o\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$. Throughout, we denote the resulting average of $o$, as a function of conserved densities $\underline{\mathsf{q}}$, by using the “sans-serif” font for the observable,
@@
for a large, macroscopic scale $\ell$. At initial time, this state is a local-equilibrium state,
$\lim_{\ell\to\infty}\langle o(\ell x,0)\rangle_{\ell}=\langle o\rangle_{\beta(x)}.$ (13)
It also has short range correlations: for every fixed $x_{1},{\rm y}_{1},x_{2},{\rm y}_{2}\in{\mathbb{R}}$, one has
$\langle o_{1}(\ell x_{1}+{\rm y}_{1},0),o_{2}(\ell x_{2}+{\rm y}_{2},0)\rangle_{\ell}^{\rm c}=\Big{(}\langle o_{1}({\rm y}_{1},0),o_{2}({\rm y}_{2},0)\rangle_{\beta(x_{1})}^{\rm c}+\mathcal{O}(\ell^{-2})\Big{)}\delta_{x_{1},x_{2}}+\mathcal{O}(\ell^{-\infty})$ (14)
where $\delta_{x_{1},x_{2}}=1$ if $x_{1}=x_{2}$ and 0 otherwise. A more precise expression of these initial correlations, as distributions on the macroscopic coordinates, is given in (24) below. At later times it develops long-range correlations [7]. More general states, with long-range correlations or with long-wavelength source insertions at various macroscopic times, can also be considered.